# Early Bright Galaxies from Helium Enhancements in High-Redshift Star Clusters

### v1

@@ Abstract
The first few cycles of JWST have identified an overabundance of UV-bright galaxies and a general excess of UV luminosity density at $z\gtrsim 10$ compared to expectations from most (pre-JWST) theoretical models. Moreover, some of the brightest high-redshift spectroscopically confirmed galaxies exhibit peculiar chemical abundance patterns, most notably extremely high N/O ratios. Since N/O has been empirically shown to scale strongly with He/H, as expected for hot hydrogen burning, these same bright high-redshift galaxies are likely also helium-enhanced. Under simplistic assumptions for stellar evolution, the bolometric luminosity of a star scales as $L\propto(4-\frac{9}{2}Y+\frac{5}{4}Y^{2})^{-1}$ — hence a higher He/H leads to brighter stars. In this Letter, we evolve a series of MESA models to the zero-age main-sequence and highlight that the helium enhancements at the levels measured and inferred for high-redshift galaxies can boost the 1500 Å UV luminosity by up to $\sim 50\%$, while simultaneously increasing the stellar effective temperature. The combination of helium enhancements with nebular continuum emission expected for intense bursts of star formation have the potential to help reduce the tension between JWST observations and certain galaxy formation models.
@@ 1 Introduction
Photometric and spectroscopic measurements from JWST (Gardner et al., 2006) have demonstrated that there are both more UV-bright galaxies at high redshift and more generally an excess UV luminosity density compared to expectations from most pre-JWST theoretical models of high-redshift galaxy formation (e.g., Finkelstein et al., 2023; Harikane et al., 2024b; Leung et al., 2023; Chemerynska et al., 2023, although cf. Willott et al. 2023). Numerous astrophysical solutions have been proposed to explain these observations that can be categorized as either (1) modifications to the observed mass-to-light ratio or (2) appealing to a higher star formation efficiency.
@@ 2 The Impact of Helium Enhancements on 1500 Å UV Luminosities
$L\propto\frac{1}{(4-\frac{9}{2}Y+\frac{5}{4}Y^{2})}.$ (2)
@@
In Figure 1 (green line) we show the bolometric luminosity of a fixed-mass star, subject to the above-listed assumptions, scaled to the luminosity assuming Big Bang Nucleosynthesis (BBN) abundances ($Y=0.24$, Planck Collaboration et al. 2020). For the maximum helium abundance seen in Omega Centauri (Dupree & Avrett, 2013) and observed in the $z\sim 6$ galaxy RXCJ2248-ID (Topping et al., 2024b; Yanagisawa et al., 2024), the bolometric luminosity increases by a factor of $\sim 50\%$.
@@
Finally, it is important to highlight that the impact on the total bolometric luminosity is not constant with mass. This is demonstrated in Figure 3, where we show the bolometric luminosity of individual stars at the ZAMS as a function of mass for different values of $Y$, normalized to the values at the same mass for $Y=0.24$, assuming a metallicity of $0.1Z_{\odot}$. We observe that lower-mass stars exhibit larger increases in bolometric luminosity for all values of $Y$ compared to higher-mass stars. This effect can be understood from homology relations. The highest mass stars have non-negligible contributions from radiation pressure to their equation of state. For a star entirely dominated by radiation pressure, at fixed mass, $L\propto(2-Y)^{-1}$. At masses $\lesssim 2\ {\rm M_{\odot}}$, free-free and bound-free absorption contribute to the opacity where it can be shown that $L\propto(2-\frac{5}{4}Y)^{-7.5}$. Because the impact of helium enhancements is mass-dependent, the evolution of properties such as mass-to-light ratio and when spectral features such as the Balmer break appear as a function of time may change and complicate estimates of stellar mass or star formation histories. These again depend on detailed calculations of stellar atmospheres and main-sequence lifetimes, but the effects may be relevant for interpreting the light from early galaxies impacted by globular cluster populations and other densely clustered star formation at a broad range of ages, in addition to the very young UV N-emitters that we have focused on in this work.
@@ 3 Discussion & Conclusions
Here we have highlighted the fact that helium enhancements, which empirically scale strongly with N/O (Yanagisawa et al., 2024), lead to a boost in UV luminosity of individual stars due to the strong scaling of luminosity with mean molecular weight. For the nitrogen emitters observed at $z>10$, detailed stellar evolution models predict that the 1500 Å UV luminosity can increase as much as $\sim 50\%$. This effect is likely to take place only in compact star-forming regions undergoing extreme bursts that exhibit similar conditions to what is expected for globular cluster ‘second population’ formation. Such galaxies represent $\sim 50\%$ of the very UV-bright population at high redshift, while the others are potentially mergers (e.g., Harikane et al., 2024a). Thus the combination of an intense burst of star formation along with the luminosity increase from an enhanced helium abundance and strong nebular continuum could help explain why there exists a population of compact, very UV-bright galaxies at high redshift. Moreover, if a significant fraction of high-redshift star formation occurs in massive star clusters as predicted from Milky Way data (Belokurov & Kravtsov, 2023), helium enhancements may be more common at high-redshift than in the local Universe; and could contribute to the general excess UV luminosity density as well as impact retrieval of stellar masses and ages for affected systems.

### v3

@@ Abstract
The first few cycles of JWST have identified an overabundance of UV-bright galaxies and a general excess of UV luminosity density at $z\gtrsim 10$ compared to expectations from most (pre-JWST) theoretical models. Moreover, some of the brightest high-redshift spectroscopically confirmed galaxies exhibit peculiar chemical abundance patterns, most notably extremely high N/O ratios. Since N/O has been empirically shown to scale strongly with He/H, as expected for hot hydrogen burning, these same bright high-redshift galaxies are likely also helium-enhanced. Under simplistic assumptions for stellar evolution, the bolometric luminosity of a star scales as $L\propto(2-\frac{5}{4}Y)^{-4}(2-Y)^{-1}$ — hence a higher He/H leads to brighter stars. In this Letter, we evolve a series of MESA models to the zero-age main-sequence and highlight that the helium enhancements at the levels measured and inferred for high-redshift galaxies can boost the 1500 Å UV luminosity by up to $\sim 50\%$, while simultaneously increasing the stellar effective temperature. The combination of helium enhancements with nebular continuum emission expected for intense bursts of star formation have the potential to help reduce the tension between JWST observations and certain galaxy formation models.
@@ 1 Introduction
Photometric and spectroscopic measurements from JWST (Gardner et al., 2006) have demonstrated that there are both more UV-bright galaxies at high redshift and more generally an excess UV luminosity density compared to expectations from most pre-JWST theoretical models of high-redshift galaxy formation (e.g., Finkelstein et al., 2023; Harikane et al., 2024; Leung et al., 2023; Chemerynska et al., 2023, although cf. Willott et al. 2023). Numerous astrophysical solutions have been proposed to explain these observations that can be categorized as either (1) modifications to the observed mass-to-light ratio or (2) appealing to a higher star formation efficiency.
@@ 2 The Impact of Helium Enhancements on 1500 Å UV Luminosities
$L\propto\frac{1}{(2-\frac{5}{4}Y)^{4}(2-Y)}.$ (2)
@@
In Figure 1 (green line) we show the bolometric luminosity of a fixed-mass star, subject to the above-listed assumptions, scaled to the luminosity assuming Big Bang Nucleosynthesis (BBN) abundances ($Y=0.24$, Planck Collaboration et al. 2020). For the maximum helium abundance seen in Omega Centauri (Dupree & Avrett, 2013) and observed in the $z\sim 6$ galaxy RXCJ2248-ID (Topping et al., 2024b; Yanagisawa et al., 2024), the bolometric luminosity increases by a factor of $\sim 80\%$.
@@
Finally, it is important to highlight that the impact on the total bolometric luminosity is not constant with mass. This is demonstrated in Figure 3, where we show the bolometric luminosity of individual stars at the ZAMS as a function of mass for different values of $Y$, normalized to the values at the same mass for $Y=0.24$, assuming a metallicity of $0.1Z_{\odot}$. We observe that lower-mass stars exhibit larger increases in bolometric luminosity for all values of $Y$ compared to higher-mass stars. This effect can be understood from homology relations. The highest mass stars have non-negligible contributions from radiation pressure to their equation of state. For a star entirely dominated by radiation pressure, at fixed mass, $L\propto(2-Y)^{-1}$. At masses $\lesssim 2\ {\rm M_{\odot}}$, free-free and bound-free absorption contribute to the opacity where it can be shown that $L\propto\mu^{7.5}$. Because the impact of helium enhancements is mass-dependent, the evolution of properties such as mass-to-light ratio and when spectral features such as the Balmer break appear as a function of time may change and complicate estimates of stellar mass or star formation histories. These again depend on detailed calculations of stellar atmospheres and main-sequence lifetimes, but the effects may be relevant for interpreting the light from early galaxies impacted by globular cluster populations and other densely clustered star formation at a broad range of ages, in addition to the very young UV N-emitters that we have focused on in this work.
@@ 3 Discussion & Conclusions
Here we have highlighted the fact that helium enhancements, which empirically scale strongly with N/O (Yanagisawa et al., 2024), lead to a boost in UV luminosity of individual stars due to the strong scaling of luminosity with mean molecular weight. For the nitrogen emitters observed at $z>10$, detailed stellar evolution models predict that the 1500 Å UV luminosity can increase as much as $\sim 50\%$. This effect is likely to take place only in compact star-forming regions undergoing extreme bursts that exhibit similar conditions to what is expected for globular cluster ‘second population’ formation. Such galaxies represent $\sim 50\%$ of the very UV-bright population at high redshift, while the others are potentially mergers (e.g., Harikane et al., 2024). Thus the combination of an intense burst of star formation along with the luminosity increase from an enhanced helium abundance and strong nebular continuum could help explain why there exists a population of compact, very UV-bright galaxies at high redshift. Moreover, if a significant fraction of high-redshift star formation occurs in massive star clusters as predicted from Milky Way data (Belokurov & Kravtsov, 2023), helium enhancements may be more common at high-redshift than in the local Universe; and could contribute to the general excess UV luminosity density as well as impact retrieval of stellar masses and ages for affected systems.